OpenStudy (anonymous):

Solve 3x - 5y = 17 9x - 15y = -4

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

what is there 2 solve? u put the answer. :/

OpenStudy (anonymous):

The answer chocies are (-1, -4) (4, -1) No Solutions Infinitely Many Solutions

OpenStudy (anonymous):

ok,

OpenStudy (anonymous):

No solutions becuz you have two variables that are still in there, you cant solve it without knowing the value of the adverb, the best u could do is simplify.

OpenStudy (anonymous):

and cant simplify becuz the variables are not the same

OpenStudy (anonymous):

So the answer would be no solutions right? thanks!

OpenStudy (anonymous):

yes :) np

OpenStudy (anonymous):

okay sorry about this but i have another problem lol its Solve x + 3y = 9 3x - 3y = -13 and the answer choices are (-1,10/3 ) (1,-10/3 ) (-1,3/10 ) (1,-3/10 )

OpenStudy (anonymous):

ok,

OpenStudy (anonymous):

why is there answers already there?

OpenStudy (anonymous):

is that the exact problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hmm :/ i have my try hard face on

OpenStudy (anonymous):

the lesson is called solving systems of equations algebraically

OpenStudy (anonymous):

if thats any help haha

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

im srry, idk this

OpenStudy (anonymous):

texaschic? can you help me lol

OpenStudy (texaschic101):

x + 3y = 9 --> x = -3y + 9 now sub -3y + 9 in for x in the other equation 3x - 3y = -13 3(-3y + 9) - 3y = -13 -9y + 27 - 3y = -13 -9y - 3y = -13 - 27 -12y = -40 y = 40/12 y = 10/3 now sub 10/3 in for y x + 3y = 9 x + 3(10/3) = 9 x + 10 = 9 x = 9 - 10 x = -1 answer is (-1,10/3)